OpenStudy (anonymous):

Express x^5/3 in simplest radical form form

Nnesha (nnesha):

same like last one try it!!

OpenStudy (anonymous):

x5sqrt(x^2)

OpenStudy (anonymous):

x^5

OpenStudy (jdoe0001):

\(\large \bf { a^{\frac{{\color{blue} n}}{{\color{red} m}}} \implies \sqrt[{\color{red} m}]{a^{\color{blue} n}} \qquad \qquad \sqrt[{\color{red} m}]{a^{\color{blue} n}}\implies a^{\frac{{\color{blue} n}}{{\color{red} m}}}\qquad thus \\ \quad \\ x^{\frac{{\color{blue} 5}}{{\color{red}3}}}\implies ? }\)

OpenStudy (anonymous):

yea so i'm right?

OpenStudy (jdoe0001):

so solve it then, as Nnesha said

OpenStudy (anonymous):

I just said the answer didn't I?

Nnesha (nnesha):

nope.....

OpenStudy (anonymous):

5sqrt(x)^3?

Nnesha (nnesha):

how did you get that ?? explain i know those 2 are ur answer choices :(

OpenStudy (anonymous):

cause of the formula you all gave me

OpenStudy (anonymous):

and that is the only one that makes since

Nnesha (nnesha):